OpenStudy (anonymous):

2p - 6(2p + 2) + 2(5p - 5) - 3(3p - 5) A) -84p - 8 B) -9p - 7 C) -84p - 7 D) -9p - 8

OpenStudy (anonymous):

You have to expand first, so: 2p - 6(2p + 2) + 2(5p - 5) - 3(3p - 5) = 2p - 6*2p + 6*2 + 2*5p - 2*5 - 3*3p + 3*5

OpenStudy (anonymous):

Okay, so would it be 2p-12p+12+10p-10-9p+15

OpenStudy (anonymous):

Yep. Then you add the ones with a 'p' in it. And the ones without. 2p-12p+12+10p-10-9p+15 = (2p-12p+10p-9p)+(12-10+15)

OpenStudy (anonymous):

-9p + 17

OpenStudy (anonymous):

Oh, I wrote wrong, Im sorry: 2p - 6*2p - 6*2 + 2*5p - 2*5 - 3*3p + 3*5= 2p-12p-12+10p-10-9p+15 (2p-12p+10p-9p)+(-12-10+15)

OpenStudy (anonymous):

-9p - 16

OpenStudy (anonymous):

wait somethings not right..

OpenStudy (anonymous):

-12-10+15 = -22+15 = -7